• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sarty

    Log entry time 07:31:04 on June 3,2000

    Entry number 42937

    keyword=shift summary

    Production running at thcm=155 (phi=0):
    ...pointing and trigger eff. checks done at end of last shift
    ...from end of previous shift:
    ........................................................................................................................tot triggers:
    .
    ...........23:39: Run 1701: run at 50 muA - S/N not good....................................0.91 M
    ...........23:50: Run 1702: same as last run - must reduce I to improve S/N......1.25 M
    .
    ...this shift:
    00:04: Run 1703: short run at 25 muA - S/N ok (see halog 42912)..................0.18 M
    00:07: Run 1704: prod run at 25 muA ................................................................1.25 M
    00:48: Run 1705: ..................................................................................................1.04 M
    01:28: Run 1706: ..................................................................................................1.25 M
    01:55: Run 1707: ..................................................................................................1.25 M
    02:32: Run 1708: ..................................................................................................1.25 M
    03:06: Run 1709: ..................................................................................................1.25 M
    04:20: Run 1710: ..................................................................................................1.25 M
    04:22: Run 1711: ..................................................................................................1.25 M
    05:00: Run 1712: ..................................................................................................1.00 M
    ...............down for 10 minutes for quick acc checks and QE measurement......
    ...............this 10 minutes lasted for 51 minutes - ended up doing some "cleaning up"
    ...............to try and improve some periodic energy fluctuations which they have been
    ...............seeing for the past day or so.
    06:20: Run 1713: ..................................................................................................1.25 M
    07:00: Run 1714: ..................................................................................................0.60 M
    ................run ended after 20 minutes due to HRSE Q3 quench fault: needs access
    ................to reset.
    .
    Total triggers so far for thcm=+155 : 13.1 M